# Composition of Scalar Mesons and Their Effects on Nuclear Matter Properties in an Extended Linear Sigma Model

### v2

@@ Abstract
It has been argued that the iso-scalar and iso-vector mesons play significant roles in nuclear matter and neutron star structures. We improve the extended linear sigma model with baryons, proposed in our previous work, by introducing the flavor structures constructed from antisymmetric tensors of chiral representations to study these physics. The parameter space of this model is refined with well-reproduced nuclear matter properties at saturation density by the lowest order Lagrangian, ensuring consistency with vacuum results, such as $f_{\pi}\approx 134\,\text{MeV}$. The anticipated plateau-like behaviors of the symmetry energy are predicted at intermediate densities, which is crucial for the consistency of GW170817 and the neutron skin thickness of $\text{Pb}^{208}$. Subsequently, neutron star structures are calculated using several parameter sets, and the results for the nuclear matter properties at saturation density align with empirical values. It is found that the neutron star structures are sensitive to the couplings between the iso-vector $a_{0}$ meson and nucleons and the four-vector meson couplings: small values of both are favorable. Meanwhile, nuclear matter properties at saturation density favor larger values of the latter and are not sensitive to the former. This signifies the statistical significance of neutron star observations when obtaining realistic chiral effective field theories or models at various densities. The parameter set favored by neutron star observations also aligns the behavior of the sound velocity with the conformal limit at high densities relevant to cores of massive stars. It is hoped that the results of this work can guide future studies on the relationship between the microscopic symmetry of strong interactions and macroscopic phenomena.
@@ I Introduction
Nuclear matter (NM), a system dominated by the nonperturbative quantum chromodynamics (QCD), has been studied for several decades. But there are still great uncertainties about its properties, especially at compact star densities, due to our little knowledge on the phase structure of strong interaction matter Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Chen (2017); Baym et al. (2018); Gil et al. (2019); Ma and Rho (2020); Li et al. (2019); Sorensen et al. (2024). The properties of NM are highly sensitive to details of nucleon interactions. Therefore, a reliable model of NM is not only significant for revealing neutron star (NS) structures Lattimer and Prakash (2001); Özel and Freire (2016), but also can shed a light on characters of the hadrons involved.
@@
In modeling NM, based on the phenomenological one-boson exchange for nucleon interactions Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989), it has long been recognized that the scalar meson resonances, such as $\sigma,f_{0}$ and $a_{0}(\delta)$, play indispensable roles by using the mean field approach (MFA) Walecka (1974); Serot and Walecka (1986, 1997) [^1] Hereafter, following the Particle Data Group, we denote the iso-vector scalar mesons as $a_{0}$. Without the iso-scalar scalar mesons $\sigma/f_{0}$, which provide the attractive force between nucleons, it’s hard to obtain the empirical values of the saturation density $n_{0}\simeq 0.16$ fm-3 and binding energy $E_{0}\simeq 16.0$ MeV Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). And the self-interactions of $\sigma$ meson is necessary to reduce the nuclear incompressibility coefficient of symmetric nuclear matter Boguta and Bodmer (1977) and other empirical values Motohiro et al. (2015). In addition to the iso-scalar scalar mesons, the iso-vector scalar mesons are important for describing the asymmetric nuclear matter and hence the properties of neutron stars Kubis and Kutschera (1997); Hofmann et al. (2001); Liu et al. (2002); Chen et al. (2007); Roca-Maza et al. (2011); Wang et al. (2014); Typel and Alvear Terrero (2020); Kong et al. (2025). Furthermore, it is found that the scalar-iso-scalar scalar meson couplings, e.g., $\sigma a_{0}^{2}$ and $\sigma^{2}a_{0}^{2}$ have large effects on both the symmetry energy $E_{\rm sym}$ and its slope $L$ Zabari et al. (2019); Kubis et al. (2020); Miyatsu et al. (2022); Li et al. (2022). Explicitly, they soften $E_{\rm sym}$ at intermediate densities but stiffen $E_{\rm sym}$ at high densities.
@@
Normally, in the models of NM, the scalar mesons are included with respect to that quantum numbers, such as parity, spin and isospin, less attention has been paid to other properties related to QCD. From QCD point of view, the structures of scalar mesons are still under debate, especially the light ones (see, e.g., Close and Tornqvist (2002) for a review). From the mass ordering of the scalar mesons below $1$ GeV, it is difficult to regard them as the pure two-quark states but an admixture of the two- and four-quark configurations. Considering that the two quark and four quark components of scalar mesons have different axial transformation properties, we recently proposed an extended linear sigma model (ELSM) with baryons, taken as diquark-quark configurations, (bELSM) to study the properties of NM Ma and Ma (2024).
@@
In bELSM, the masses of the mesons and baryons are generated from the spontaneous chiral symmetry breaking. Due to the conservation of axial transformation, only the 2-quark component of the scalar mesons couples to the baryons at the lowest order, which may provide deeper insights into the constituents of the scalar mesons from the NM and NS properties. The $\rm U(1)_{A}$ anomaly is involved in the ’t Hooft determinant Rosenzweig et al. (1980); Kawarabayashi and Ohta (1980); Hsu et al. (1998); Fariborz et al. (2005). Another important feature of the bELSM for the phenomenological analysis is the inclusion of all possible types of mesons below $1\ \rm GeV$ and hyperons which significantly affect NS structures Gal et al. (2016).
@@
After pinning down the NM properties at saturation density $n_{0}$, a plateau-like structure of the symmetry energy, $E_{\rm sym}$, at intermediate densities, $\sim 2n_{0}$, was found in the bELSM. This is consistent with the analysis based on various of approaches in Refs. Paeng et al. (2017); Ma and Rho (2019); Ma et al. (2019); Ma and Rho (2021); Lee et al. (2022); Zabari et al. (2019); Miyatsu et al. (2022); Li et al. (2022), and satisfies the constraints of $\rm{}^{208}Pb$ Adhikari et al. (2021); Reed et al. (2021) and tidal deformations of NS from GW170817 Abbott et al. (2016). We also investigated the effects of parameter variations on different coefficients of Taylor expansions of energy per-nucleon, $E=\mathcal{E}/n$, by density, $n=n_{n}+n_{p}$, and asymmetry, $\alpha=(n_{n}-n_{p})/n$, to provide further guidance on the interaction parameterization of NM:
$E(n,\alpha)$ $\simeq$ $E(n_{0},0)+E_{\rm sym}(n_{0})\alpha^{2}+L(n_{0})\alpha^{2}\chi+\frac{K(n_{0})}{2!}\chi^{2}+\frac{J(n_{0})}{3!}\chi^{3}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)$ (1)
$\simeq$ $E_{0}+\left.\frac{\alpha^{2}}{2}\frac{\partial^{2}E(n_{0},\alpha)}{\partial\alpha^{2}}\right|_{\alpha=0}+\alpha^{2}\chi\left.\left(3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}\right)\right|_{n=n_{0}}+\frac{\chi^{2}}{2!}\left.\left(9n^{2}\frac{\partial^{2}E_{0}(n)}{\partial n^{2}}\right)\right|_{n=n_{0}}$
${}+\frac{\chi^{3}}{3!}\left.\left(27n^{3}\frac{\partial^{3}E_{0}(n)}{\partial n^{3}}\right)\right|_{n=n_{0}}+\mathcal{O}\left(\alpha^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left(\alpha^{2}\chi^{2}\right)\ ,$
@@
However, when calculating the $E_{\rm sym}(n)$ using bELSM, the plateau-like structure is found to be lower than the results in Ref. Miyatsu et al. (2022); Li et al. (2022), leading to a loose structure of NS with tidal deformation $\Lambda_{1.4}>1000$. This is due to the couplings related to $\rho$ and $\omega$ mesons being constrained by the $\rm U(3)_{V}$ symmetry, which is not entirely realistic, especially in nuclear medium Paeng et al. (2012). Therefore, improving the bELSM to resolve these problems and achieve a better description of the NM properties and NS structures is the purpose of this work.
@@
When constructing the interactions, especially in the baryonic sector, only operators like ${\rm Tr}(\bar{B}\gamma_{\mu}V^{\mu}B)$ were considered in Ref. Ma and Ma (2024), which make the couplings $g_{\rho NN}$ and $g_{\omega NN}$ the same. However, another type of operator, $\epsilon_{abc}\epsilon^{def}\bar{B}_{ad}\gamma_{\mu}B_{be}V^{\mu}_{cf}$ Papazoglou et al. (1998), which can split the couplings of the singlet and octet vector mesons to baryons. In this work, we extend the bELSM by including this kind of operators in both the mesonic and baryonic sectors at the lowest order and improve the parameter space of the original bELSM. The 12-quark terms, which were necessary to maintain reasonable behavior at high densities in the work of Ref. Ma and Ma (2024), can be removed, enhancing the rationality of the power-counting in the bELSM. The vacuum expectation values (VEVs) of the scalar meson fields are close to the results of Ref. Fariborz et al. (2008), leading to a reasonable $f_{\pi}$ (quark condensate) and low-energy pion scattering results.
@@
With the improved bELSM at the lowest order, the NM properties at $n_{0}$ and the bare masses of the mesons and baryons can be well reproduced. The bELSM is then applied to calculating NS structures. It is found that NS structures are sensitive to the NM properties at $n_{0}$, especially the incompressibility, $K(n_{0})$, which also affects the sound velocity (SV) behavior at various densities. The relationship between NS and SV is very subtle, as detailed in Refs. Bedaque and Steiner (2015); Zhang et al. (2024a). When $K(n_{0})$ falls within empirical constraints, the SV softens significantly, causing the maximum mass of NS to be lower than the observed value, $2.14\pm 0.10M_{\odot}$ Cromartie et al. (2019). To meet the constraints of the maximum mass of NS and the mass-radius (M-R) relation of NS from GW170817 Abbott et al. (2016), $K(n_{0})$ should be larger than the empirical constraints, approximately $500\ \rm MeV$. This $K(n_{0})$ also results in an interesting behavior of SV, approaching the conformal limit at high densities, $n\sim 5n_{0}$, whereas other SVs from smaller $K(n_{0})$ will be too soft. This highlights the statistical significance of NS observations when obtaining realistic chiral effective field theories ($\chi$EFTs) or models Guo et al. (2024a). In addition, the influence of portions in the 2-quark and 4-quark constituents of the scalar mesons on NS structures and NM properties is also shown and their physical magnitudes are estimated. These discussions setup a relationship between microscopic symmetry and macroscopic phenomena.
@@
The rest of this article is organized as follows: In Sec. II, we first introduce the improved bELSM at the lowest order under the relativistic mean field (RMF) approximation. Then we study the properties of the scalar mesons on the NM properties and NS structures in detail. In Sec. III, the discussion on the results of the current work and the outlook are provided. The details of the model construction and approximation are given in Appendix.
II Extended linear sigma model at densities
@@ II.1 Theoretical framework
In this work, we consider the lowest-order bELSM given in Ref. Ma and Ma (2024) including terms up to 8-quark lines with a single trace and operators of dimension-4. In addition, as mentioned in the Introduction, we consider the terms with three-rank Levi-Civita tensor in flavor space. The Lagrangian contributing under RMF approximation is given by
@@
The $h,\ e$ and $a_{1}$ terms are the terms which were not considered in Ref. Ma and Ma (2024). The details of the model construction are given in App. A.
@@
with $\lambda_{3}$ and $\lambda_{8}$ being the 3rd and 8th components of Gell-Mann matrices, and $I$ as the identity matrix. The similar decomposition applies to the four-quark configuration $\hat{S}^{\prime}$. $V$ is the zero component of vector fields which survives in the MFA
@@
So, with $p$ and $n$, there are iso-scalar scalar configurations $\sigma^{\prime}/\hat{\sigma}^{\prime}$, $f_{0}^{\prime}/\hat{f}_{0}^{\prime}$ and iso-vector scalar configurations $a_{0}^{\prime}/\hat{a}_{0}^{\prime}$ in model (2). After chiral symmetry breaking, the mixing of these scalar configurations yields the physical states of scalar mesons. The spontaneous chiral symmetry breaking is achieved by the vacuum expectation values of both 2-quark and 4-quark fields: $\langle\sigma^{\prime}\rangle=\sqrt{3}\alpha$ and $\langle\hat{\sigma}^{\prime}\rangle=\sqrt{3}\beta$, which are determined by the minimum of the potential $\mathcal{L}_{\rm M}$:
$\left\langle\frac{\partial V_{\rm M}}{\partial\sigma^{\prime}}\right\rangle$ $=$ $2\alpha\left(-c_{2}+2c_{4}\alpha^{2}+4e_{3}\beta\right)=0\ ,$
$\left\langle\frac{\partial V_{\rm M}}{\partial\hat{\sigma}^{\prime}}\right\rangle$ $=$ $2\left(d_{2}\beta+2e_{3}\alpha^{2}\right)=0\ ,$ (7)
where the three-flavor symmetry is used. The physical states $\sigma$, $a_{0}$, and $f_{0}$ are obtained by the mixing of the 2-quark and 4-quark configurations, via

### v3

@@ Abstract
It has been argued that the isoscalar and isovector mesons play significant roles in nuclear matter and neutron star structures. We improve the extended linear sigma model with baryons, proposed in our previous work, by introducing the flavor structures constructed from antisymmetric tensors of chiral representations to study these physics. The parameter space of this model is refined with well-reproduced nuclear matter properties at saturation density by the lowest-order Lagrangian, ensuring consistency with vacuum results, such as $f_{\pi}\approx 134\,\text{MeV}$. The anticipated plateaulike behaviors of the symmetry energy are predicted at intermediate densities, which is crucial for the consistency of GW170817 and the neutron skin thickness of $\text{Pb}^{208}$. Subsequently, neutron star structures are calculated using several parameter sets, and the results for the nuclear matter properties at saturation density align with empirical values. It is found that the neutron star structures are sensitive to the couplings between the isovector $a_{0}$ meson and nucleons and the four-vector meson couplings: small values of both are favorable. Meanwhile, nuclear matter properties at saturation density favor larger values of the latter and are not sensitive to the former. This signifies the statistical significance of neutron star observations when obtaining realistic chiral effective field theories or models at various densities. The parameter set favored by neutron star observations also aligns the behavior of the sound velocity with the conformal limit at high densities relevant to cores of massive stars. It is hoped that the results of this work can guide future studies on the relationship between the microscopic symmetry of strong interactions and macroscopic phenomena.
@@ I Introduction
Nuclear matter (NM), a system dominated by the nonperturbative QCD, has been studied for several decades. But there are still great uncertainties about its properties, especially at compact star densities, due to our little knowledge on the phase structure of strong interaction matter Fukushima and Hatsuda (2011); Lattimer and Prakash (2016); Chen (2017); Baym et al. (2018); Gil et al. (2019); Ma and Rho (2020); Li et al. (2019); Sorensen et al. (2024). The properties of NM are highly sensitive to details of nucleon interactions. Therefore, a reliable model of NM is not only significant for revealing neutron star (NS) structures Lattimer and Prakash (2001); Özel and Freire (2016), but also can shed a light on characters of the hadrons involved.
@@
In modeling NM, based on the phenomenological one-boson exchange for nucleon interactions Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989), it has long been recognized that the scalar meson resonances, such as $\sigma,f_{0}$ and $a_{0}$, play indispensable roles by using the mean field approach (MFA) Walecka (1974); Serot and Walecka (1986, 1997) [^1] Hereafter, following the Particle Data Group, we denote the isovector scalar mesons as $a_{0}$. Without the isoscalar scalar mesons $\sigma/f_{0}$, which provide the attractive force between nucleons, it’s hard to obtain the empirical values of the saturation density $n_{0}\simeq 0.16$ fm-3 and binding energy $E_{0}\simeq 16.0$ MeV Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989). And the self-interactions of the $\sigma$ meson are necessary to reduce the nuclear incompressibility coefficient of symmetric nuclear matter Boguta and Bodmer (1977) and other empirical values Motohiro et al. (2015). In addition to the isoscalar scalar mesons, the isovector scalar mesons are important for describing the asymmetric nuclear matter and hence the properties of neutron stars Kubis and Kutschera (1997); Hofmann et al. (2001); Liu et al. (2002); Chen et al. (2007); Roca-Maza et al. (2011); Wang et al. (2014); Typel and Alvear Terrero (2020); Kong et al. (2025). Furthermore, it is found that the scalar-isoscalar scalar meson couplings, e.g., $\sigma a_{0}^{2}$ and $\sigma^{2}a_{0}^{2}$ have large effects on both the symmetry energy $E_{\rm sym}$ and its slope $L$ Zabari et al. (2019); Kubis et al. (2020); Miyatsu et al. (2022); Li et al. (2022). Explicitly, they soften $E_{\rm sym}$ at intermediate densities but stiffen $E_{\rm sym}$ at high densities.
@@
Normally, in the models of NM, the scalar mesons are included with respect to that quantum numbers, such as parity, spin and isospin, less attention has been paid to other properties related to QCD. From a QCD point of view, the structures of scalar mesons are still under debate, especially the light ones (see, e.g., Ref. Close and Tornqvist (2002) for a review). From the mass ordering of the scalar mesons below $1$ GeV, it is difficult to regard them as the pure two-quark states but an admixture of the two- and four-quark configurations. Considering that the two-quark and four-quark components of scalar mesons have different axial transformation properties, we recently proposed an extended linear sigma model (ELSM) with baryons, where baryons are taken as diquark-quark configurations, to study the properties of NM, and the model is noted as baryonic extended linear sigma model (bELSM). Ma and Ma (2024).
@@
In bELSM, the masses of the mesons and baryons are generated from the spontaneous chiral symmetry breaking. Because of the conservation of axial transformation, only the two-quark component of the scalar mesons couples to the baryons at the lowest-order, which may provide deeper insights into the constituents of the scalar mesons from the NM and NS properties. The $\rm U(1)_{A}$ anomaly is involved in the ’t Hooft determinant Rosenzweig et al. (1980); Kawarabayashi and Ohta (1980); Hsu et al. (1998); Fariborz et al. (2005). Another important feature of the bELSM for the phenomenological analysis is the inclusion of all possible types of mesons below $1\ \rm GeV$ and hyperons which significantly affect NS structures Gal et al. (2016).
@@
After pinning down the NM properties at saturation density $n_{0}$, a plateaulike structure of the symmetry energy, $E_{\rm sym}$, at intermediate densities, $\sim 2n_{0}$, was found in the bELSM. This is consistent with the analysis based on various of approaches in Refs. Paeng et al. (2017); Ma and Rho (2019); Ma et al. (2019); Ma and Rho (2021); Lee et al. (2022); Zabari et al. (2019); Miyatsu et al. (2022); Li et al. (2022), and satisfies the constraints of $\rm{}^{208}Pb$ Adhikari et al. (2021); Reed et al. (2021) and tidal deformations of NS from GW170817 Abbott et al. (2016). We also investigated the effects of parameter variations on different coefficients of Taylor expansions of energy per nucleon, $E=\mathcal{E}/n$, by density, $n=n_{n}+n_{p}$, and asymmetry, ${\delta}=(n_{n}-n_{p})/n$, to provide further guidance on the interaction parameterization of NM,
$E(n,{\delta})$ $\simeq$ $E(n_{0},0)+E_{\rm sym}(n_{0}){\delta}^{2}+L(n_{0}){\delta}^{2}\chi+\frac{K(n_{0})}{2!}\chi^{2}+\frac{J(n_{0})}{3!}\chi^{3}+\mathcal{O}\left({\delta}^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left({\delta}^{2}\chi^{2}\right)$ (1)
$\simeq$ $E_{0}+\left.\frac{{\delta}^{2}}{2}\frac{\partial^{2}E(n_{0},{\delta})}{\partial{\delta}^{2}}\right|_{{\delta}=0}+{\delta}^{2}\chi\left.\left(3n\frac{\partial E_{\mathrm{sym}}(n)}{\partial n}\right)\right|_{n=n_{0}}+\frac{\chi^{2}}{2!}\left.\left(9n^{2}\frac{\partial^{2}E_{0}(n)}{\partial n^{2}}\right)\right|_{n=n_{0}}$
${}+\frac{\chi^{3}}{3!}\left.\left(27n^{3}\frac{\partial^{3}E_{0}(n)}{\partial n^{3}}\right)\right|_{n=n_{0}}+\mathcal{O}\left({\delta}^{4}\right)+\mathcal{O}\left(\chi^{4}\right)+\mathcal{O}\left({\delta}^{2}\chi^{2}\right)\ ,$
@@
However, when calculating the $E_{\rm sym}(n)$ using bELSM, the plateaulike structure is found to be lower than the results in Refs. Miyatsu et al. (2022); Li et al. (2022), leading to a loose structure of NS with tidal deformation $\Lambda_{1.4}>1000$. This is due to the couplings related to $\rho$ and $\omega$ mesons being constrained by the $\rm U(3)_{V}$ symmetry, which is not entirely realistic, especially in the nuclear medium Paeng et al. (2012). Therefore, improving the bELSM to resolve these problems and achieve a better description of the NM properties and NS structures is the purpose of this work.
@@
When constructing the interactions, especially in the baryonic sector, only operators like ${\rm Tr}(\bar{B}\gamma_{\mu}V^{\mu}B)$ were considered in Ref. Ma and Ma (2024), which make the couplings $g_{\rho NN}$ and $g_{\omega NN}$ the same. However, another type of operator, $\epsilon_{abc}\epsilon^{def}\bar{B}_{ad}\gamma_{\mu}B_{be}V^{\mu}_{cf}$ Papazoglou et al. (1998), which can split the couplings of the singlet and octet vector mesons to baryons. In this work, we extend the bELSM by including this kind of operators in both the mesonic and baryonic sectors at the lowest-order and improve the parameter space of the original bELSM. The 1two-quark terms, which were necessary to maintain reasonable behavior at high densities in the work of Ref. Ma and Ma (2024), can be removed, enhancing the rationality of the power counting in the bELSM. The vacuum expectation values of the scalar meson fields are close to the results of Ref. Fariborz et al. (2008), leading to a reasonable $f_{\pi}$ (quark condensate) and low-energy pion scattering results.
@@
With the improved bELSM at the lowest-order, the NM properties at $n_{0}$ and the bare masses of the mesons and baryons can be well reproduced. The bELSM is then applied to calculating NS structures. It is found that NS structures are sensitive to the NM properties at $n_{0}$, especially the incompressibility, $K(n_{0})$, which also affects the sound velocity (SV) behavior at various densities. The relationship between NS and SV is very subtle, as detailed in Refs. Bedaque and Steiner (2015); Zhang et al. (2024a). When $K(n_{0})$ falls within empirical constraints, the SV softens significantly, causing the maximum mass of NS to be lower than the observed value, $2.14\pm 0.10M_{\odot}$ Cromartie et al. (2019). To meet the constraints of the maximum mass of NS and the mass-radius (M-R) relation of NS from GW170817 Abbott et al. (2016), $K(n_{0})$ should be larger than the empirical constraints, approximately $500\ \rm MeV$. This $K(n_{0})$ also results in an interesting behavior of SV, approaching the conformal limit at high densities, $n\sim 5n_{0}$, whereas other SVs from smaller $K(n_{0})$ will be too soft. This highlights the statistical significance of NS observations when obtaining realistic chiral effective field theories or models Guo et al. (2024a). In addition, the influence of portions in the two-quark and four-quark constituents of the scalar mesons on NS structures and NM properties is also shown and their physical magnitudes are estimated. These discussions set up a relationship between microscopic symmetry and macroscopic phenomena.
@@
The rest of this article is organized as follows: In Sec. II, we first introduce the improved bELSM at the lowest-order under the relativistic mean field (RMF) approximation. Then we study the properties of the scalar mesons on the NM properties and NS structures in detail. In Sec. III, the discussion on the results of the current work and the outlook are provided. The details of the model construction and approximation are given in the Appendix.
II Extended linear sigma model at finite densities
@@ II.1 Theoretical framework
In this work, we consider the lowest-order bELSM given in Ref. Ma and Ma (2024) including terms up to eight quark lines with a single trace and operators of dimension 4. In addition, as mentioned in the Introduction, we consider the terms with three-rank Levi-Civita tensor in flavor space. The Lagrangian contributing under RMF approximation is given by
@@
The h,eh,\ e and $a_{1}$ terms are the terms which were not considered in Ref. Ma and Ma (2024). The details of the model construction are given in Appendix A.
@@
with $\lambda_{3}$ and $\lambda_{8}$ being the third and eighth components of Gell-Mann matrices and $I$ being the identity matrix. The similar decomposition applies to the four-quark configuration $\hat{S}^{\prime}$. $V$ is the zero component of vector fields which survives in the MFA
@@
So, with $p$ and $n$, there are isoscalar scalar configurations $\sigma^{\prime}/\hat{\sigma}^{\prime}$, $f_{0}^{\prime}/\hat{f}_{0}^{\prime}$ and isovector scalar configurations $a_{0}^{\prime}/\hat{a}_{0}^{\prime}$ in model (2). After chiral symmetry breaking, the mixing of these scalar configurations yields the physical states of scalar mesons. The spontaneous chiral symmetry breaking is achieved by the vacuum expectation values of both two-quark and four-quark fields: $\langle\sigma^{\prime}\rangle=\sqrt{3}\alpha$ and $\langle\hat{\sigma}^{\prime}\rangle=\sqrt{3}\beta$, which are determined by the minimum of the potential $\mathcal{L}_{\rm M}$,
$\left\langle\frac{\partial\mathcal{L}_{\rm M}}{\partial\sigma^{\prime}}\right\rangle$ $=$ $2\alpha\left(-c_{2}+2c_{4}\alpha^{2}+4e_{3}\beta\right)=0\ ,$
$\left\langle\frac{\partial\mathcal{L}_{\rm M}}{\partial\hat{\sigma}^{\prime}}\right\rangle$ $=$ $2\left(d_{2}\beta+2e_{3}\alpha^{2}\right)=0\ ,$ (7)
where the three-flavor symmetry is used. The physical states $\sigma$, $a_{0}$, and $f_{0}$ are obtained by the mixing of the two-quark and four-quark configurations, via